Burrell Company purchased a machine for $43,000 on January 2, 2019. The machine has an estimated service life of 5 years and a zero estimated residual value. The asset earns income before depreciation and income taxes of $21,500 each year. The tax rate is 30%.

Required:

Compute the rate of return earned (on the average net asset value) by the company each year of the asset's life under the straight-line and the double-declining-balance depreciation methods. Assume that the machine is the company's only asset.

Straight-line method. If required, round to one decimal place.

2019 %
2020 %
2021 %
2022 %
2023 %


Double-declining-balance depreciation method. Round to two decimal places. Round your intermediate dollar value calculations to the nearest whole number.

2019 %
2020 %
2021 %
2022 %
2023 %
Answer #1
Machine cost $        43,000
Less: Residual value $                 -  
Depreciable cost $        43,000
Divided by: Estimated service life                       5
Depreciation Expense per Year $          8,600
Year Beginning asset value Less: Depreciation Expense Ending asset value
2019 $                43,000 $                 8,600 $        34,400
2020 $                34,400 $                 8,600 $        25,800
2021 $                25,800 $                 8,600 $        17,200
2022 $                17,200 $                 8,600 $          8,600
2023 $                  8,600 $                 8,600 $                 -  
Straight-line method
Year 2019 2020 2021 2022 2023
Income before depreciation and taxes $                21,500 $               21,500 $        21,500 $            21,500 $         21,500
Less: Depreciation Expense $                  8,600 $                 8,600 $          8,600 $              8,600 $            8,600
Income before taxes $                12,900 $               12,900 $        12,900 $            12,900 $         12,900
Less: Income taxes (Income before taxes * 30%) $                  3,870 $                 3,870 $          3,870 $              3,870 $            3,870
Net income $                  9,030 $                 9,030 $          9,030 $              9,030 $            9,030
Beginning asset value $                43,000 $               34,400 $        25,800 $            17,200 $            8,600
Ending asset value $                34,400 $               25,800 $        17,200 $              8,600 $                   -  
Total $                77,400 $               60,200 $        43,000 $            25,800 $            8,600
Average asset value (Total / 2) $                38,700 $               30,100 $        21,500 $            12,900 $            4,300
Net income $                  9,030 $                 9,030 $          9,030 $              9,030 $            9,030
Divided by: Average asset value $                38,700 $               30,100 $        21,500 $            12,900 $            4,300
Rate of return earned 23.3% 30.0% 42.0% 70.0% 210.0%
Depreciation rate under straight-line method (1/5) 20%
Depreciation rate under the double-declining-balance depreciation method (2*20%) 40%
Double-declining-balance depreciation method
Year Beginning asset value Less: Depreciation Expense Ending asset value Calculation for depreciation
2019 $                43,000 $               17,200 $        25,800 Depreciation = 43000 * 40% = 17200
2020 $                25,800 $               10,320 $        15,480 Depreciation = 25800 * 40% = 10320
2021 $                15,480 $                 6,192 $          9,288 Depreciation = 15480 * 40% = 6192
2022 $                  9,288 $                 3,715 $          5,573 Depreciation = 9288 * 40% = 3715
2023 $                  5,573 $                 5,573 $                 -   Depreciation = Beginning Assets value - Residual value = 5573-0 = 5573
Double-declining-balance depreciation method
Year 2019 2020 2021 2022 2023
Income before depreciation and taxes $                21,500 $               21,500 $        21,500 $            21,500 $         21,500
Less: Depreciation Expense $                17,200 $               10,320 $          6,192 $              3,715 $            5,573
Income before taxes $                  4,300 $               11,180 $        15,308 $            17,785 $         15,927
Less: Income taxes (Income before taxes * 30%) $                  1,290 $                 3,354 $          4,592 $              5,336 $            4,778
Net income $                  3,010 $                 7,826 $        10,716 $            12,449 $         11,149
Beginning asset value $                43,000 $               25,800 $        15,480 $              9,288 $            5,573
Ending asset value $                25,800 $               15,480 $          9,288 $              5,573 $                   -  
Total $                68,800 $               41,280 $        24,768 $            14,861 $            5,573
Average asset value (Total / 2) $                34,400 $               20,640 $        12,384 $              7,431 $            2,787
Net income $                  3,010 $                 7,826 $        10,716 $            12,449 $         11,149
Divided by: Average asset value $                34,400 $               20,640 $        12,384 $              7,431 $            2,787
Rate of return earned 8.75% 37.92% 86.53% 167.53% 400.04%